OpenStudy (anonymous):

2-|3x+1|=6 the answers are -5/3 and 1 When the answers, -5/3 and 1, are plugged into the problem for x the solution does not equal 6. Why?

OpenStudy (anonymous):

That Is The w

OpenStudy (anonymous):

That Is The Way It Is

OpenStudy (anonymous):

Are you sure you typed the problem correctly?

OpenStudy (amistre64):

2-|3x+1|=6 -|3x+1| = 4 ; no solution

OpenStudy (amistre64):

2 - (what positive value) = 6?